OpenStudy (anonymous):

1 MEDAL QUESTION!!!!!!! The Frostburg-Truth bus travels on a straight road from Frostburg Mall to Sojourner Truth Park. The mall is 2 miles east and 5 miles north of the City Center. The park is 5 miles west and 5 miles south of the Center. How far is it from the mall to the park to the nearest tenth of a mile? 5.4 miles 12.2 miles 13.5 miles 7.1 miles

OpenStudy (anonymous):

there is no drawing

OpenStudy (anonymous):

im just confused

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

much better but how do i get the answer

OpenStudy (anonymous):

well. a^2+b^2=c^2

OpenStudy (anonymous):

that means that leg1^2+leg2^2= hypotenuse^2

OpenStudy (anonymous):

PT

OpenStudy (anonymous):

and the square root of the hypotenuse would be the length of that rite triangle. once you do that to both triangles add both of those square roots to get your answer

OpenStudy (anonymous):

quick question is this question multiple choice?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

no A. 5.4 miles B.12.2 miles C. 13.5 miles D.7.1 miles

OpenStudy (anonymous):

well first try to solve the substitute in the equation

OpenStudy (jdoe0001):

 use pythagorean theorem as already suggested by iplayffxiv